OpenStudy (anonymous):

find all solutions for the equation (or inequality),find all real solutions of the equation. t^6+t^3-2=0

OpenStudy (lgbasallote):

let t^3 = x so t^6 = x^2 x^2 + x -2 =0 factor.. (x+2)(x-1) =0 so x= -2 x= 1 but x= t^3 so t= cuberoot of -2 and t = 1

OpenStudy (anonymous):

thank you